A cylindrical solid has a base diameter of 10cm and height of 20cm. the solid also weighs 4500g. calculate the density of cylindrical solid in g/cm3

1
Expert's answer
2021-08-26T07:16:31-0400

Diameter of cylinder = "10\space cm"

Height of cylinder = "20\space cm"

Volume of cylinder = "\dfrac{\pi d^2h}{4}=1570.79\space cm^3"

Weight of solid = 4500 g

Density = "\dfrac{Mass}{Volume}=\dfrac{4500\space g}{1570.79\space cm^3}=2.86\space g/cm^3"